20191209 UT

Like the previous night, this one, too, was to be split between SOUL ESM testing (1st half) and ARGOS (2nd half), with D/E programs filling in when ARGOS could not propagate.

The enclosure could not be opened because of heavy clouds and snow.

OSA: J Williams
Observers: G Rahmer and S Walsh
ISA: O Kuhn

Shane and Gustavo ran up ARGOS, using the ARGOS cal source and diodes for the NGS & lasers. They could close the loop, despite some struggles to get the WFS pupils aligned and the 'star' on the tip of the pyramid (a problem only seen with the closed dome setup, and not on-sky).

They also noted vibrations.

With the loop closed, the LUCI image did not look too elongated but the images were being taken using the N3.75 camera.

During the first hours HBS was off, but starting at 21:30, HBS was turned on. The HBS on/off test was done last night and had little effect on the vibrations.

-- OlgaKuhn - 09 Dec 2019
Topic revision: r1 - 09 Dec 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback